# Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting

### v1

@@ A.2 Training VLAs
A.3 Generating Textual Descriptions
We require short and long textual descriptions of each personal object to construct hard and short prompt baselines. To obtain these descriptions, we use GPT-4o with following prompts (for long and short each). For each object, we choose one canonical reference view (front-facing if available) and query the model once.
Prompt Used to Generate Long Textual Description for Hard Prompt Methods Describe only the object in the image. Focus on its most distinctive features such as shape, color, and material. Write the result in English as a short noun phrase, not a full sentence. Do not mention orientation, position, or surroundings.
Prompt Used to Generate Short Textual Description for Hard Prompt Methods Describe only the object in the image. Write the result in English as a single, highly detailed noun phrase, not a full sentence. Include as many intrinsic and identifying features as possible, such as overall shape, dimensions, proportions, material, surface finish, textures, patterns, colors, edges, rims, openings, and any decorative or structural details. Do not mention orientation, position, or surroundings.
We manually verify that the generated phrases correctly describe each object and lightly edit obviously incorrect attributes when necessary. We then use the short descriptions as hard prompts and the long descriptions as rich prompts for our personalization baselines. Table LABEL:tab:textual_descriptions in the appendix lists the final short and long descriptions for objects in Personalized-SIMPLER, Personalized-VLABench, and the real-world scenarios.
@@ A.3 Generating Textual Descriptions
marker red and white cylindrical board marker with ridged cap a cylindrical board marker with a white plastic body featuring bold red text and graphics, including the words M̈YPEN Board Markeränd a stylized pen icon, accented by red bands near the ends, a ribbed red cap, and a smooth, glossy surface finish
@@ Appendix B Token Learning Strategy
In addition to VAP, we consider a token-based personalization strategy as a natural competitor, inspired by recent work on personalizing vision–language models and image generation models. In this line of work, a new token embedding is learned to represent a user-specific concept, while keeping the backbone model frozen. Methods such as MyVLM (alaluf2024myvlm) and Yo’LLaVA (nguyen2024yollavapersonalizedlanguagevision) follow this paradigm and have shown strong personalization performance in question answering by learning concept tokens with carefully constructed hard and easy negatives.
@@ B.1 Learning Token Embeddings on the Backbone VLM
We closely follow the training pipeline of Yo’LLaVA (nguyen2024yollavapersonalizedlanguagevision) when learning personalized tokens in the VLM. Concretely, we start from a frozen PaliGemma 3B model (beyer2024paligemma), which combines a SigLIP image encoder with a Gemma-based language decoder and is designed for transfer to downstream vision–language tasks. For each personal object, we introduce a new subject identifier token (e.g., <my_cup>) into the vocabulary together with a small number of learnable soft tokens. The rest of the VLM parameters, including the vision encoder and language decoder, remain fixed during training.
@@
Positive examples use the user-provided reference images of $o$. For hard negatives, instead of mining from a large external dataset as in Yo’LLaVA (nguyen2024yollavapersonalizedlanguagevision), we leverage the actual distractors that appear in our personalized benchmarks. For each distractor instance, we capture more than ten views and label them as hard negatives for the corresponding object token. Easy negatives are sampled from objects that never appear as distractors for $o$ (e.g., props from other tasks), providing a diverse background of unrelated images. This choice is arguably favorable to the token-learning baseline, since the hard negatives are drawn directly from the deployment distribution; nevertheless, our method VAP still substantially outperforms this baseline in personalized manipulation.
@@ B.2 Token Embedding Space Analysis
Our proxy strategy relies on the assumption that the token embedding space of the VLM remains largely unchanged when the model is fine-tuned into a VLA. To verify this, we compare the token embeddings of PaliGemma 3B (used as our VLM) with those of the corresponding VLA ($\pi_{0}$ and $\pi_{0.5}$), which is obtained by training on large-scale robotic action data (black2025pi05). Both models share the same tokenizer and vocabulary, so their embeddings are aligned by construction.
We extract the token embedding matrices from the text decoders of PaliGemma 3B and each VLA variant, and evaluate their alignment using three metrics: token-wise cosine similarity, linear centered kernel alignment (CKA), and a k-nearest-neighbor top-1 (kNN@1) token matching accuracy. CKA measures how similar the pairwise inner-product structure of two representation spaces is, and equals $1$ when one space can be obtained from the other by an orthogonal transformation. The kNN@1 score measures, for each VLM token, whether its nearestz neighbor in the VLA embedding space is the corresponding token, thus directly quantifying how often token identities are preserved.
@@
These observations are consistent with prior findings on language model fine-tuning, which report that lower layers and embedding matrices change little compared to upper layers during task-specific adaptation (merchant2020happens; hao2020investigating). In our context, they suggest that learning personalized tokens on the PaliGemma VLM and then transferring them to the VLA is a reasonable approximation: the token embeddings live in essentially the same space before and after action fine-tuning. This motivates our use of VLM-trained concept tokens as a strong token-learning baseline for personalized manipulation.
@@ B.3 Attention Analysis
Concretely, we extract the VLA’s cross-attention weights for the query position corresponding to the learned object token, and take the attention mass over image-token keys as a per-patch score from the last layer (layer 17). We then reshape and resize these scores to the image resolution and normalize them across spatial locations to obtain a heatmap. Intuitively, this heatmap measures how strongly the concept token “recognizes” each region of the image. Figure 9B.3 shows representative examples from Personalized-SIMPLER, overlaid on the original RGB observations.
@@
One plausible explanation for the remaining failures—including episodes where the heatmap appears relatively consistent on the target yet the behavior is still incorrect—is the VLA architecture. The vision–language encoder that produces the text/image embeddings (and contains the learned concept token) is shared with the underlying VLM, whereas the action head is a separate expert trained via flow matching. During action prediction, latent action variables are sampled from noise and attend over the entire sequence of text and image embeddings through cross-attention, mixing information across modalities and tokens. As a result, even when the learned token induces a stable preference in the perceptual representations, this fine-grained, instance-specific cue can be diluted or overridden before it consistently affects the final low-level action distribution. This suggests that, beyond improving token-induced localization, bridging token-level conditioning into reliable instance-aware control in a separate action expert remains challenging.
@@ D.2 Case 2: Wrong Visual Prompt (Multi-View Inconsistency)
The second category arises from inconsistencies across camera views and accounts for a large fraction of failures in multi-view settings (59.2% in Personalized-VLABench and 50.9% in Real-world; Table 9). Even when each individual view is moderately informative, per-view matching and tracking can occasionally lock onto different objects in different cameras, especially under severe occlusion, partial visibility in the wrist camera, or extreme viewpoint changes that alter apparent shape/texture. This produces multi-view scenes where one camera correctly highlights the personal object, while another highlights a distractor, yielding conflicting prompts that can destabilize behavior or bias actions toward the wrong view. These cases expose a limitation of our current per-view grounding design and motivate future work on explicitly enforcing cross-view consistency. For example, recent cross-view correspondence/segmentation work such as ObjectRelator introduces mechanisms for cross-view object alignment across ego-/exo-centric perspectives that may provide useful building blocks for reducing Case 2 failures in our setting (fu2025objectrelator) (see Figure 19).
@@ D.3 Case 3: Correct Visual Prompt but Failed Manipulation
Appendix E Discussion and Future Directions

### v2

@@ A.3 Instruction Rewriting Logic
A.4 Generating Textual Descriptions
We require short and long textual descriptions of each personal object to construct hard and short prompt baselines. To obtain these descriptions, we use GPT-5.1 with the following prompts (for long and short each). For each object, we use all reference views and query the model once.
Prompt Used to Generate Short Textual Description for Hard Prompt Methods Describe only the object in the image. Focus on its most distinctive features such as shape, color, and material. Write the result in English as a short noun phrase, not a full sentence. Do not mention orientation, position, or surroundings.
Prompt Used to Generate Long Textual Description for Hard Prompt Methods Describe only the object in the image. Write the result in English as a single, highly detailed noun phrase, not a full sentence. Include as many intrinsic and identifying features as possible, such as overall shape, dimensions, proportions, material, surface finish, textures, patterns, colors, edges, rims, openings, and any decorative or structural details. Do not mention orientation, position, or surroundings.
We use the short descriptions as hard prompts and the long descriptions as rich prompts for our personalization baselines. Table LABEL:tab:textual_descriptions in the appendix lists the final short and long descriptions for objects in Personalized-SIMPLER, Personalized-VLABench, and the real-world scenarios.
@@ A.4 Generating Textual Descriptions
pouch small round beige fabric pouch with zipper and matching strap a small soft-sided light gray rounded zippered pouch with a slightly flattened oval profile, finely ribbed synthetic fabric front panel, thick padded binding around the perimeter, continuous zipper track along the rim with braided-edge seam detail, gold-toned metal zipper hardware, and an attached matching woven webbing wrist strap loop with reinforced stitching and metal ring connector
@@ Appendix B Token Learning Strategy
In addition to VAP, we consider a token-based personalization strategy as a natural competitor, inspired by recent work on personalizing vision–language models and image generation models. In this line of work, a new token embedding is learned to represent a user-specific concept, while keeping the backbone model frozen. Methods such as MyVLM (Alaluf et al., 2024) and Yo’LLaVA (Nguyen et al., 2024) follow this paradigm and have shown strong personalization performance in question answering by learning concept tokens with carefully constructed hard and easy negatives.
@@ B.1 Learning Token Embeddings on the Backbone VLM
We closely follow the training pipeline of Yo’LLaVA (Nguyen et al., 2024) when learning personalized tokens in the VLM. Concretely, we start from a frozen PaliGemma 3B model (Beyer et al., 2024), which combines a SigLIP image encoder with a Gemma-based language decoder and is designed for transfer to downstream vision–language tasks. For each personal object, we introduce a new subject identifier token (e.g., <my_cup>) into the vocabulary together with a small number of learnable soft tokens. The rest of the VLM parameters, including the vision encoder and language decoder, remain fixed during training.
@@
Positive examples use the user-provided reference images of $o$. For hard negatives, instead of mining from a large external dataset as in Yo’LLaVA (Nguyen et al., 2024), we leverage the actual distractors that appear in our personalized benchmarks. For each distractor instance, we capture more than ten views and label them as hard negatives for the corresponding object token. Easy negatives are sampled from objects that never appear as distractors for $o$ (e.g., props from other tasks), providing a diverse background of unrelated images. This choice is favorable to the token-learning baseline and can be viewed as a form of evaluation-set leakage, since the hard negatives are drawn directly from the deployment-time distractor instances. We adopt this setting intentionally to strengthen the baseline; thus, the improvements of VAP over Soft Prompt should be interpreted as conservative.
@@ B.2 Token Embedding Space Analysis
Our proxy strategy relies on the assumption that the token embedding space of the VLM remains largely unchanged when the model is fine-tuned into a VLA. To verify this, we compare the token embeddings of PaliGemma 3B (used as our VLM) with those of the corresponding VLA ($\pi_{0}$ and $\pi_{0.5}$), which is obtained by training on large-scale robotic action data (Intelligence et al., 2025). Both models share the same tokenizer and vocabulary, so their embeddings are aligned by construction.
We extract the token embedding matrices from the text decoders of PaliGemma 3B and each VLA variant, and evaluate their alignment using three metrics: token-wise cosine similarity, linear centered kernel alignment (CKA), and a k-nearest-neighbor top-1 (kNN@1) token matching accuracy. CKA measures how similar the pairwise inner-product structure of two representation spaces is, and equals $1$ when one space can be obtained from the other by an orthogonal transformation. The kNN@1 score measures, for each VLM token, whether its nearest neighbor in the VLA embedding space is the corresponding token, thus directly quantifying how often token identities are preserved.
@@
These observations are consistent with prior findings on language model fine-tuning, which report that lower layers and embedding matrices change little compared to upper layers during task-specific adaptation (Merchant et al., 2020; Hao et al., 2020). In our context, they suggest that learning personalized tokens on the PaliGemma VLM and then transferring them to the VLA is a reasonable approximation: the token embeddings live in essentially the same space before and after action fine-tuning. This motivates our use of VLM-trained concept tokens as a strong token-learning baseline for personalized manipulation.
@@ B.3 Attention Analysis
Concretely, we extract the VLA’s cross-attention weights for the query position corresponding to the learned object token, and take the attention mass over image-token keys as a per-patch score from the last layer (layer 17). We then reshape and resize these scores to the image resolution and normalize them across spatial locations to obtain a heatmap. Intuitively, this heatmap measures how strongly the concept token “recognizes” each region of the image. Figure 9B.3 show representative examples from Personalized-SIMPLER, overlaid on the original RGB observations.
@@
One plausible explanation for the remaining failures—including episodes where the heatmap appears relatively consistent on the target yet the behavior is still incorrect—is the VLA architecture. The vision–language encoder that produces the text/image embeddings (and contains the learned concept token) is shared with the underlying VLM, whereas the action head is a separate expert trained via flow matching. During action prediction, latent action variables are sampled from noise and attend over the entire sequence of text and image embeddings through cross-attention, mixing information across modalities and tokens. As a result, even when the learned token induces a stable preference in the perceptual representations, this fine-grained, instance-specific cue can be diluted or overridden before it consistently affects the final low-level action distribution. This suggests that, beyond improving token-induced localization, bridging token-level conditioning into reliable instance-aware control in a separate action expert remains challenging. While our analysis focuses on the $\pi$-family backbone used in this paper, similar modular designs with a language/vision trunk and a separate action head are common in contemporary VLAs, so this limitation may extend beyond this specific backbone.
@@ Appendix C Qualitative Examples
Figure 18: Qualitative Examples on Real-world Scenario - 1
Figure 19: Qualitative Examples on Real-world Scenario - 2
Figure 20: Qualitative Examples on Real-world Scenario - 3
@@ D.2 Case 2: Wrong Visual Prompt (Multi-View Inconsistency)
The second category arises from inconsistencies across camera views and accounts for a large fraction of failures in multi-view settings (59.2% in Personalized-VLABench and 50.9% in Real-world; Table 9). Our current system factorizes grounding and tracking per camera view for real-time inference: retrieval/matching and mask propagation are performed independently in each view, and the resulting per-view masks are directly used as prompts. Under severe occlusion, partial visibility (especially in the wrist camera), or extreme viewpoint changes, this factorized design can cause per-view matching/tracking to lock onto different same-category instances across cameras. Consequently, one camera may correctly highlight the personal object while another highlights a distractor, yielding contradictory prompts that can destabilize behavior or bias actions toward the wrong view/object.
Note that the voting mechanism (Sec. 3.3) aggregates evidence across reference images within each view to select the target proposal. However, it does not couple camera views or enforce a shared cross-view identity during grounding, and thus cannot prevent per-view drift under occlusion or eliminate contradictory prompts across cameras. These cases expose a limitation of our current independent per-view design and motivate future work on explicitly enforcing cross-view consistency.
Future direction: cross-view association for joint evidence fusion (MAP-style).
Case 2 fundamentally stems from the lack of an explicit shared target identity across views: our current pipeline performs retrieval/matching and tracking independently per camera, so there is no latent instance variable that couples observations across cameras. A natural extension is therefore to introduce an intermediate cross-view association stage that enforces a single physical instance identity across views before generating per-view prompts, and then aggregates multi-view evidence while keeping the downstream VLA policy frozen.
Formally, let each view $v\in\{1,\dots,V\}$ produce candidate instance proposals $\mathcal{B}^{(v)}=\{b^{(v)}_{i}\}_{i=1}^{M_{v}}$ (e.g., detector masks/boxes with short-horizon tracks) with proposal embeddings $e^{(v)}_{i}$ (already computed for reference matching), and let the reference set $R_{o}$ provide embeddings $\{z_{r}\}_{r\in R_{o}}$. We represent the shared identity by a per-view assignment vector $m=\{m_{v}\}_{v=1}^{V}$ where $m_{v}\in\{1,\dots,M_{v}\}\cup\{\varnothing\}$ selects which proposal corresponds to the target in view $v$ (or $\varnothing$ when the target is unobserved due to occlusion). A MAP-style discrete approximation is to solve
$m^{*}=\arg\max_{m}\;\;\underbrace{\sum_{v=1}^{V}\phi_{\text{ref}}\!\left(v,m_{v}\right)}_{\text{reference evidence}}\;+\;\lambda\underbrace{\sum_{v<u}\phi_{\text{cv}}\!\left(v,m_{v};\,u,m_{u}\right)}_{\text{cross-view consistency}}\;+\;\beta\underbrace{\sum_{v=1}^{V}\phi_{\text{obs}}(v,m_{v})}_{\text{visibility/track prior}},$
where $\phi_{\text{ref}}$ measures reference agreement (e.g., $\phi_{\text{ref}}(v,i)=\sum_{r\in R_{o}}\cos(e^{(v)}_{i},z_{r})$ and $\phi_{\text{ref}}(v,\varnothing)=0$), $\phi_{\text{obs}}$ encodes optional unary priors (e.g., detector confidence or track stability), and $\phi_{\text{cv}}$ encourages that selected proposals across views correspond to the same physical instance. In the following, we keep the unary terms $\phi_{\text{ref}}$ and $\phi_{\text{obs}}$ fixed and discuss alternative instantiations/approximations for $\phi_{\text{cv}}$ and the resulting joint inference. Below we outline a few plausible design directions for cross-view association, rather than an exhaustive treatment.
Embedding-based clustering. A simple instantiation sets $\phi_{\text{cv}}(v,i;\,u,j)=\cos(e^{(v)}_{i},e^{(u)}_{j})$ (and 0 if either side is $\varnothing$), reusing already-computed proposal embeddings. A common approximation is to build a cross-view graph over proposals with edge weights given by $\phi_{\text{cv}}$, cluster proposals into cross-view instance hypotheses $\mathcal{K}$, and then select the cluster maximizing aggregated unary evidence, $\mathrm{score}(k)=\sum_{(v,i)\in k}\big(\phi_{\text{ref}}(v,i)+\beta\,\phi_{\text{obs}}(v,i)\big)$. Per-view prompts are then produced from members of the chosen cluster, enforcing agreement by construction.
Bipartite matching and track consistency. Another standard approximation solves pairwise assignments between view pairs. For a pair $(v,u)$, one can define a cost that combines appearance and optional geometry (when available), e.g.,
$\mathrm{cost}(i,j)=-\cos(e^{(v)}_{i},e^{(u)}_{j})+\gamma\,d_{\mathrm{geom}}\!\left(b^{(v)}_{i},b^{(u)}_{j}\right)$,
and apply Hungarian matching to obtain consistent correspondences, which can then be merged into multi-view tracklets/clusters followed by the same multi-view evidence aggregation.
Geometry-aware coupling when calibration is available. When camera calibration and robot kinematics are available, geometric consistency can further instantiate $\phi_{\text{cv}}$ via reprojection constraints. For example, letting $c^{(v)}_{i}$ denote a proposal centroid/keypoints in view $v$ and $\Pi_{u}(\cdot)$ the projection into view $u$, one can use a coarse 3D proxy $\hat{X}_{i}$ (e.g., triangulation from multiple views, a depth prior, or kinematic constraints) and define
$\phi_{\text{cv}}(v,i;\,u,j)=\cos(e^{(v)}_{i},e^{(u)}_{j})-\eta\left\|\Pi_{u}\!\left(\hat{X}_{i}\right)-c^{(u)}_{j}\right\|_{2}\quad(\text{or IoU between a projected region and }b^{(u)}_{j})$.
This formalizes the intuition that calibration/kinematics can couple views through reprojection, but it is sensitive to calibration/synchronization and the quality of the 3D proxy and may introduce new failure modes if applied naively.
Learned cross-view correspondence as a building block. Alternatively, learned cross-view correspondence/segmentation modules (e.g., ObjectRelator) may provide building blocks to associate per-view proposals across ego-/exo-centric perspectives and enforce agreement (Fu et al., 2025). In the above formulations, such modules can instantiate $\phi_{\text{cv}}$ via learned association scores (or a soft correspondence field), which can be particularly helpful when geometric cues are weak or unavailable.
Why this is non-trivial (expected challenges). While the objective is conceptually clean, naively adding cross-view association can introduce a new bottleneck. Cross-view matching is itself error-prone under heavy occlusion and extreme viewpoint changes, and erroneous associations can be worse than leaving views independent by forcing all prompts to agree on an incorrect identity. Moreover, proposal embeddings may not be sufficiently view-invariant for near-identical instances, clustering/assignment can fragment a true instance or merge multiple instances, and geometry-based coupling depends on calibration accuracy and reliable 3D proxies. Finally, joint inference increases computation and debugging complexity, requiring careful approximations and uncertainty handling to maintain real-time performance. We therefore leave explicit cross-view association and joint evidence fusion as important directions for future work.
@@ D.3 Case 3: Correct Visual Prompt but Failed Manipulation
Appendix E Ablation Studies
This section ablates key design choices in Visual Attentive Prompting (VAP) to clarify which components are responsible for the gains reported in Section 5. We conduct all ablations on Personalized-SIMPLER to isolate the effects of grounding and prompt design under a controlled visual setting. Multi-view benchmarks introduce an additional confound—cross-view association and consistency—that can dominate failure cases (Section 5.5). We therefore analyze multi-view inconsistency separately and focus here on component-level causality.
Unless otherwise stated, all variants use the same frozen VLA policy and perception backbones as VAP, and follow the same evaluation protocol and metrics (Section 5). Tables 1013 summarize the results.
E.1 Rewriting the Instruction: Is It Necessary?
VAP combines a visual highlight with a lightweight instruction rewrite (Section 3.3) to align the language referent with the prompted region. To test whether rewriting is a critical ingredient, we compare three variants: (i) VAP (mask + rewrite), (ii) mask-only (visual highlight with the original instruction unchanged), and (iii) rewrite-only (rewritten instruction without any visual highlight). Table 10 reports performance on Google Robot and WidowX in Personalized-SIMPLER.
Results and Analysis.
Table 10 shows that combining the visual highlight with instruction rewriting is essential for strong performance. VAP achieves 62.7% SR / 83.7% CMR on Google Robot and 83.5% SR / 93.6% CMR on WidowX, whereas removing rewriting (Mask-only) collapses performance to 7.3%/20.4% (SR/CMR) and 5.2%/23.9%, respectively. This indicates that a visual marker alone is not reliably exploited by the frozen VLA when the instruction remains in a personalized form (e.g., “my X”). Instead, language–vision alignment via rewriting is a critical ingredient. Conversely, rewriting without any visual prompt (Rewrite-only) provides only limited gains. Note that our rewrite template explicitly introduces the tint-color word (e.g., “the red cc”) even when the highlight is absent; thus, Rewrite-only serves as a negative control that tests whether a color-based linguistic cue alone can drive personalization in a cluttered scene. The limited performance indicates that the gains of VAP do not come from language rewriting in isolation, but from the coupled language–vision alignment created by both the mask highlight and the rewrite.
Method Personalized-SIMPLER (Google Robot) Personalized-SIMPLER (WidowX)
CMR (%) $\uparrow$ SR (%) $\uparrow$ CMR (%) $\uparrow$ SR (%) $\uparrow$
VAP (mask + rewrite) 83.7 62.7 93.6 83.5
Mask-only 20.4 7.3 23.9 5.2
Rewrite-only 35.2 15.8 41.7 18.8
Table 10: Ablation of instruction rewriting on single-view Personalized-SIMPLER. “Mask-only” removes rewriting, while “Rewrite-only” removes the visual highlight but keeps the same tint-color rewrite template, serving as a negative control for language-only cueing.
E.2 Aggregating Evidence Across References: Voting vs. Averaging
In the grounding function $g$, VAP selects the target proposal by aggregating reference-view similarity via a voting rule (Section 3.3). A natural alternative is to average similarity scores across references and pick the maximum. We compare these two aggregation schemes while keeping all other components fixed. Table 11 summarizes the results.
Results and Analysis.
Table 11 compares two aggregation rules for selecting the target proposal across reference images. Voting yields consistently stronger results than averaging, with modest gains on Google Robot (83.7 vs. 82.0 CMR; 62.7 vs. 62.3 SR) and more pronounced gains on WidowX (93.6 vs. 87.4 CMR; 83.5 vs. 80.6 SR). While both rules perform competitively, the consistent improvement motivates our default choice. Beyond accuracy, voting provides an interpretable consensus across reference views: each reference contributes a discrete top-1 preference, and the final selection is the proposal supported by the majority, which aligns naturally with our non-parametric reference-memory framing and simplifies diagnosis of grounding decisions.
Aggregation Personalized-SIMPLER (Google Robot) Personalized-SIMPLER (WidowX)
CMR (%) $\uparrow$ SR (%) $\uparrow$ CMR (%) $\uparrow$ SR (%) $\uparrow$
Voting (VAP) 83.7 62.7 93.6 83.5
Averaging 82.0 62.3 87.4 80.6
Table 11: Ablation of reference aggregation on single-view Personalized-SIMPLER. “Voting” is the default rule in VAP (Section 3.3), while “Averaging” selects proposals by mean cosine similarity across references.
E.3 Sensitivity to the Number of Reference Images
VAP uses a small set of reference images as a non-parametric memory (Section 3.1). To quantify sensitivity to reference availability, we subsample the reference pool and evaluate VAP with $K\in\{1,3,5,7\}$ reference images.
Results and Analysis.
Table 12 shows a clear benefit from increasing the number of reference images. With a single reference (K=1K{=}1), VAP already performs reasonably on both robots, indicating a practical low-reference regime. Increasing to K=3K{=}3 yields the largest improvement, consistent with the value of additional viewpoint coverage for distinguishing near-identical instances. Moving from K=3K{=}3 to K=5K{=}5 provides further gains, while increasing to K=7K{=}7 yields only marginal additional improvement, suggesting that performance begins to saturate with a small handful of references. Based on this trade-off, we use K=5K{=}5 as the default setting throughout the paper, capturing most of the benefit while keeping the reference collection burden and initialization cost modest.
$K$ Personalized-SIMPLER (Google Robot) Personalized-SIMPLER (WidowX)
CMR (%) $\uparrow$ SR (%) $\uparrow$ CMR (%) $\uparrow$ SR (%) $\uparrow$
$K=1$ 60.2 45.6 73.7 61.9
$K=3$ 81.2 59.4 90.2 78.4
$K=5$ 83.7 62.7 93.6 83.5
$K=7$ 84.1 62.8 93.1 85.2
Table 12: Sensitivity to the number of reference images $K$ on single-view Personalized-SIMPLER. We subsample the reference set and report SR/CMR for Google Robot and WidowX.
E.4 Visual Prompt Design: Mask vs. Bounding Box and Highlight Strength
Finally, we ablate the visual prompt design used to inject grounding into the frozen VLA (Section 3.3). We evaluate two aspects: (i) the prompt geometry, comparing a segmentation-mask tint (default in VAP) against a bounding-box-based prompt, and (ii) the highlight strength, varying the overlay opacity $\alpha$. Table 13 summarizes both factors.
Results and Analysis.
Table 13 highlights that both the geometry and strength of the visual prompt matter. First, segmentation-mask highlighting consistently outperforms bounding-box prompting at the same opacity, suggesting that instance-aligned prompts provide a cleaner and less ambiguous attention signal for manipulation. In contrast, bounding boxes often include background and nearby distractors, which can dilute the conditioning and interfere with fine-grained action execution. Second, the opacity sweep reveals a saliency–occlusion trade-off: very weak highlights are insufficient to reliably guide the policy, while overly strong overlays can obscure visual cues relevant for precise grasping and placement. Overall, intermediate opacity values perform best, motivating our default choice ($\alpha{=}0.5$) used in VAP.
Prompt Variant Personalized-SIMPLER (Google Robot) Personalized-SIMPLER (WidowX)
CMR (%) $\uparrow$ SR (%) $\uparrow$ CMR (%) $\uparrow$ SR (%) $\uparrow$
Mask (VAP, $\alpha=0.5$) 83.7 62.7 93.6 83.5
Box ($\alpha=0.5$) 54.0 33.9 65.6 47.2
Mask ($\alpha=0.1$) 43.2 22.5 47.9 26.0
Mask ($\alpha=0.3$) 75.2 52.4 80.5 73.2
Mask (VAP, $\alpha=0.5$) 83.7 62.7 93.6 83.5
Mask ($\alpha=0.7$) 80.1 61.5 94.2 83.2
Table 13: Ablation of visual prompt design on single-view Personalized-SIMPLER. We compare segmentation-mask highlighting against bounding-box prompts, and sweep the overlay opacity $\alpha$. Here the default opacity is $\alpha=0.5$.
Appendix F Practical Usage: Multi-user Personal Objects in a Shared Scene
In practical deployments, multiple users often share a workspace and issue personalized requests that refer to different users’ belongings (e.g., “my brother’s dog figurine” vs. “my ornament”). While VAP is primarily an input-side personalization mechanism, it naturally extends to such shared-scene settings. Our key design choice is to execute multi-object requests sequentially, applying a single mask prompt at a time, which avoids overwhelming the frozen VLA with multiple concurrent visual cues. This section provides a small real-world demonstration of this capability; we do not claim that the VLA infers ownership from vision alone, but rather that VAP correctly binds ownership terms to visual prompts and executes the resulting sub-goals reliably.
Setup.
We use the real-world SO-101 tabletop setup (Section 4.2). The scene contains two user-associated target objects—a brother-dog figurine and an ornament—along with 2 distractor objects from the same categories for each target. For each target, we collect K=5K{=}5 reference images from the real cameras to form the corresponding reference set $R_{o}$.
Unified Sequential Execution Strategy.
We consider a compound instruction involving both targets, e.g., “put my brother’s dog figurine into the plastic bowl, and then put my ornament into the plastic bowl.” We assume a lightweight high-level planner (e.g., rule-based decomposition or an off-the-shelf LLM) to convert the request into two ordered sub-goals. The system then executes these sub-goals sequentially to maximize reliability. For each sub-goal, it retrieves the corresponding reference set, applies the standard VAP pipeline to generate a single target mask (and corresponding instruction rewrite), and executes the action with the frozen VLA. After completing the first action, the mask is cleared and VAP is re-run on the updated scene to re-ground the second target, making the second action robust to state changes caused by the first manipulation.
Metrics.
Each trial consists of two sub-tasks (brother’s dog figurine, then ornament). We report: (i) 2-step Success Rate (both sub-tasks succeed), (ii) Wrong-Object Rate (the robot manipulates a distractor object), and (iii) Others (all remaining failures, e.g., grasp/placement failures or collisions despite targeting the intended object).
Results.
As shown in Table 14, the unified sequential strategy achieves a $65.0\%$ 2-step success rate; among failures, $20.0\%$ are due to manipulating distractors and $15.0\%$ fall into Others. Overall, these results suggest that VAP can be readily used in shared-scene, multi-user settings by decomposing multi-object requests into sequential sub-goals with single-target prompting. In practice, this provides a simple and reliable path to extending VAP beyond single-object personalization without modifying the frozen VLA.
Method 2-step SR (%) $\uparrow$ Wrong-obj. (%) $\downarrow$ Others (%) $\downarrow$
Unified Sequential (Single-mask VAP) 65.0 20.0 15.0
Table 14: Multi-user practical usage in a shared scene. We execute a two-object request sequentially with a single target mask at a time. Reported over 20 randomized trials with 2 distractor objects from the same categories for each target object.
Qualitative Examples.
Figure 24 visualizes representative rollouts, overlaying the VAP mask at each sub-task to make the grounding decision explicit. The figure highlights how the system focuses on one target at a time (brother’s dog figurine first, then ornament) and updates the mask after the first action changes the scene.
Figure 24: Unified sequential execution with a single mask. VAP grounds and manipulates my brother’s dog figurine. After the first action perturbs the scene, VAP is re-run to correctly locate and mask my ornament for the second sub-task, illustrating robust binding under state changes.
Appendix G Discussion and Future Directions